OpenStudy (anonymous):

A list of 5 positive integers has the following properties: the only integer in the list that occurs more than once is 8; the median of the numbers in the list is 9; and the average of the numbers in the list is 10. What is the largest possible integer that could appear in the list?

OpenStudy (amistre64):

Its been awhile, but does median mean the middle number, Ex: (1,4,9) median=4?

OpenStudy (amistre64):

and does the list have to go from smallest to largest?

OpenStudy (amistre64):

I think I am right; so lets try this: (8,8,9,x,y) and (8+8+9+x+y)/5 = 10 25+x+y= 50 x+y=25 10 is the next smallest interger; and 10 + 15 = 25. So I would say that 15 is the largest interger that could apper in the list.

OpenStudy (amistre64):

the more math I learn, the worse my spelling gets :)

OpenStudy (anonymous):

that looks right thank you soo much